# Sparse Autoencoder Neural Operators: Model Recovery in Function Spaces

### v2

@@ Abstract
We frame the problem of unifying representations in neural models as one of sparse model recovery and introduce a framework that extends sparse autoencoders (SAEs) to lifted spaces and infinite-dimensional function spaces, enabling mechanistic interpretability of large neural operators (NO). While the Platonic Representation Hypothesis suggests that neural networks converge to similar representations across architectures, the representational properties of neural operators remain underexplored despite their growing importance in scientific computing. We compare the inference and training dynamics of SAEs, lifted-SAE, and SAE neural operators. We highlight how lifting and operator modules introduce beneficial inductive biases, enabling faster recovery, improved recovery of smooth concepts, and robust inference across varying resolutions, a property unique to neural operators.
@@ 1 Introduction
A growing body of work reports that neural networks, despite architectural differences, converge to unifying representations across contexts such as brain–model comparisons [1, 2], alignment with large language models [3], brain–computer interfaces [4], and universality in dynamics [5]. To study this, methods for representation similarity have been developed [6, 7, 8], building on representational similarity analysis [9]. The Platonic Representation Hypothesis (PRH) [10] formalizes why and where such convergence arises, framing it as a unifying principle across architectures. Recent work suggests it may arise from implicit regularization in stochastic gradient descent, which favours solutions with minimal gradient norms [11]. These advances raise a central question: under what conditions do different architectures learn equivalent representations, and how can this be formalized?
@@
If convergence is truly universal, it should also extend beyond Euclidean spaces to neural operators (NO) [12, 13, 14], which learn mappings between infinite-dimensional function spaces. Neural operators are now central to data-driven scientific modelling [15] and have recently been applied in neuroscience to capture brain representations [16, 17]. While their predictive power is well established across applications [18, 19, 20], their representational properties and interpretability remain largely unknown, leaving open whether universal representation principles extend naturally to function spaces.
To study the PRH in this broader setting, we adopt the perspective of concept learning [21], focusing on sparse model recovery [22, 23]. Rather than asking whether two models learn the same representations, we ask whether they recover the same underlying concepts, an approach aligned with mechanistic interpretability, where SAEs extract interpretable features from large networks [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. By framing representation learning as estimating latent sparse codes and dictionaries, we unify comparisons across architectures, training dynamics, and inductive biases. This leads us to three central questions: i) Can recovery generalize from Euclidean to functional spaces? ii) Under what conditions do networks and operators recover equivalent representations, and when do operators offer advantages? iii) How does lifting affect recovery dynamics?
Our Contributions We address these questions by extending SAEs to lifted SAEs (L-SAEs), which learn concepts in a lifted space, and SAE neural operators (SAE-NOs), which formalize model recovery in function spaces. We show that lifting enables recovery while acting as a preconditioner that accelerates learning, and identify conditions where the inference and dynamics of L-SAEs reduce to those of standard SAEs. We further introduce SAE-NOs, demonstrating that i) lifting-induced preconditioning extends to L-SAE-NOs, ii) truncated modes confer an inductive bias that is advantageous for recovering smooth concepts, iii) SAE-NOs recover representations robustly under resolution changes where SAEs fail, and iv) with full-frequency filters and matched spatial support, their dynamics reduce to those of SAE-CNN.
Fig. 1: Model Recovery with SAEs. a) Architectural comparison of SAE, lifted SAE, and SAE Neural Operators. b) Learning in sampled Euclidean spaces vs. function spaces.
@@
Preliminaries The sparse linear generative model (Def.˜1.1), also known as sparse coding [22, 23], has a rich history in inverse problems [35], deeply studied in statistics [36] and in compressed sensing [37, 38, 39, 40]. Representing data $\bm{x}$} under a sparse generative model involves learning a dictionary $\bm{D}$$^{*}$ to infer a sparse code $\bm{z}$}. We focus on the case where $\bm{D}$$^{*}$ is unknown (a.k.a. dictionary learning[41, 42, 43] and refer to it as sparse model recovery, formulated as a bilevel problem (Def.˜1.2[44].
@@
A data sample ${\bm{x}}\in\mathbb{R}^{m}$ is said to follow a sparse generative model if there exists a sparse latent representation ${\bm{z}}\in\mathbb{R}^{p}$ (with $\text{supp}({\bm{z}})\leq k\ll p$) and an overcomplete dictionary ${\bm{D}}^{*}\in\mathbb{R}^{m\times p}$ (with $p\gg m$) that model the data as ${\bm{x}}={\bm{D}}^{*}{\bm{z}}$ in the noiseless setting (see Def.˜B.1 for sparse convolutional generative models).
Definition 1.2 (Sparse Model Recovery).
@@
Under the sparse generative model (Def.˜1.1), the goal of sparse model recovery is to estimate the underlying dictionary $\bm{D}$$^{*}$ by solving the following problem:
$\text{(outer)}\ \min_{{\bm{D}}\in\mathcal{D}}\quad\tfrac{1}{2}\ \|{\bm{x}}-{\bm{D}}{\bm{z}}\|_{2}^{2}+\lambda\mathcal{R}({\bm{z}})\quad\text{s.t.}\quad\text{(inner)}\ {\bm{z}}=f_{\theta}({\bm{x}})=\operatorname*{arg\,min}_{{\bm{v}}\in\mathbb{R}^{p}}\mathcal{L}({\bm{x}},{\bm{v}},\theta),$ (1)
@@
where $f_{\theta}$ is a forward map from data $\bm{x}$} to latent code $\bm{z}$}, defined as the minimizer of the inner loss $\mathcal{L}({\bm{x}},{\bm{v}},\theta)$. The outer-level optimizes $\bm{D}$}, and model recovery is successful when the learned concept $\bm{D}$} approximates the true underlying dictionary $\bm{D}$$^{*}$.
@@
The connection between bilevel optimization and neural networks is well established through SAEs [45, 46, 47, 48]. The inner and outer levels in Def.˜1.2 mirror the encoder–decoder structure of a neural network [44], where the encoder $f_{\theta}$ maps data $\bm{x}$} to sparse codes $\bm{z}$}, and the decoder reconstructs $\bm{x}$} via ${\bm{D}}{\bm{z}}$ using learned weights $\bm{D}$} (see Def.˜1.3).
Definition 1.3 (Sparse Autoencoders for Model Recovery).
@@
The goal of model recovery with SAEs is to infer a sparse code ${\bm{z}}\in\mathbb{R}^{p}$, representing the data ${\bm{x}}\in\mathbb{R}^{m}$, with a linear combination of a dictionary $\bm{D}$}: $\quad\text{(encoder)}\quad{\bm{z}}=f_{\theta}({\bm{x}}),\quad\text{(decoder)}\quad\hat{\bm{x}}={\bm{D}}{\bm{z}}$, where $\theta$ denotes the set of encoder parameters. Model recovery refers to learning the dictionary $\bm{D}$} (see Def.˜B.1 for SAE-CNN).
Sparse model recovery has attracted interest from two communities: a) unrolled learning, using optimization-inspired architectures for network design [46, 47, 48, 49, 50, 51, 52, 53, 54]; and b) sparse interpretability or mechanistic interpretability, employing SAEs to extract concepts from large models [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. To extend SAEs to function spaces, we first cover basics of neural operators [12, 13, 14]. Unlike neural networks, which operate on finite-dimensional Euclidean spaces, neural operators learn mappings between infinite-dimensional function spaces over bounded domains. They have gained traction in data-driven scientific modelling [15], particularly for solving partial differential equations (PDEs), and can generalize across discretizations. Neural operators consist of the following three modules of lifting, kernel integration, and projection (see Appendix and [14]).
@@
Fourier Neural Operator (FNO) [13] models the kernel integral operator with a convolution operator parameterized in Fourier space (Def.˜1.4, see also Def.˜B.3 for Fourier transform). The convolution operator parameterizes the kernel $\kappa(x,y)=\kappa(x-y)$ as a complex function $\kappa:D\rightarrow\mathbb{C}^{d_{w}\times d_{v}}$ (Def.˜B.2). FNO is used with truncated frequency modes in practice; this has been shown to improve performance and lower sensitivity to change (decreasing) the discretization sampling [13, 14].
Definition 1.4 (Fourier integral operator $\mathcal{K}$ (restated from [13, 14])).
Define the Fourier integral operator by $(\mathcal{K}(\phi){\bm{v}})({\bm{x}})=\mathcal{F}^{-1}\big(R_{\phi}\cdot(\mathcal{F}{\bm{v}})\big)({\bm{x}}),\ \forall{\bm{x}}\in D$, where $R_{\phi}\coloneqq\mathcal{F}(\kappa)$ is the Fourier transform of a periodic function $\kappa:D\rightarrow\mathbb{C}^{d_{w}\times d_{v}}$ parameterized by $\phi$.
To extend SAEs to function spaces, we leverage the link between convolution in Euclidean space and multiplication in the Fourier domain. Building on [55] showing SAE-induced inductive biases, we investigate when networks and operators recover similar or distinct concepts. Section˜2 formalizes sparse model recovery with neural operators [12, 13, 14] for concept learning in function spaces.
2 SAE Neural Operators for Model Recovery in Function Spaces
We extend SAEs, which are typically implemented with MLP layers (SAE-MLPs), to neural operators to enable concept recovery in function spaces. Based on operator learning, we formulate a sparse functional generative model of data (Def.˜2.1), and its corresponding optimization for model recovery on function spaces (Def.˜2.2).
Definition 2.1 (Sparse Functional Generative Model).
@@ 2 SAE Neural Operators for Model Recovery in Function Spaces
Consider a linear operator $\mathcal{G}_{{\bm{D}}^{*}}:\mathcal{Z}\rightarrow\mathcal{X}$ in function spaces, where $\mathcal{Z}$ and $\mathcal{X}$ are Banach spaces of functions defined on bounded domains $D_{z}\subset\mathbb{R}^{p}$ and $D_{x}\subset\mathbb{R}^{m}$, respectively. The data sample ${\bm{x}}\in\mathcal{X}$ is said to follow a sparse functional generative model if there exists a sparse latent representation $\bm{z}$} drawn from the support $\mathcal{Z}$ of a probability measure $\mu$, where ${\bm{x}}=\mathcal{G}_{{\bm{D}}^{*}}({\bm{z}})$.
Definition 2.2 (Sparse Functional Model Recovery).
@@
Under the sparse functional generative model (Def.˜2.1), the goal of sparse model recovery in function spaces is to estimate the underlying operator $\mathcal{G}_{{\bm{D}}^{*}}$ by solving the following bilevel optimization problem:
$\min_{{\bm{D}}\in\mathcal{D}}\quad\ \mathbb{E}_{{\bm{z}}\sim\mu}\|\mathcal{G}_{{\bm{D}}^{*}}({\bm{z}})-\mathcal{G}_{{\bm{D}}}(\hat{\bm{z}})\|_{\mathcal{X}}^{2}\quad\text{s.t.}\quad\hat{\bm{z}}=\operatorname*{arg\,min}_{{\bm{v}}}\ \mathbb{E}_{{\bm{z}}\sim\mu}\|{\bm{v}}-\mathcal{H}_{\theta}(\mathcal{G}_{{\bm{D}}^{*}}({\bm{z}}))\|_{\mathcal{Z}}^{2},$ (2)
@@
where the operator $\mathcal{H}_{\theta}:\mathcal{X}\rightarrow\mathcal{Z}$ aims to approximate the inverse of the operator $\mathcal{G}_{{\bm{D}}^{*}}$. Model recovery is successful when the learned concepts $\bm{D}$} in $\mathcal{G}_{{\bm{D}}}$ closely approximate $\bm{D}$$^{*}$ in $\mathcal{G}_{{\bm{D}}^{*}}$.
@@
To recover concepts, we formalize SAE-NO (Def.˜2.3) with an encoder approximating the inverse generative operator and a decoder that captures the underlying operator and its learnable concepts.
Definition 2.3 (Sparse Autoencoder Neural Operators (SAE-NOs) for Model Recovery).

### v3

# Mechanistic Interpretability with Sparse Autoencoder Neural Operators

@@ Abstract
We introduce sparse autoencoder neural operators (SAE-NOs), a new class of sparse autoencoders that operate directly in infinite-dimensional function spaces. We generalize the linear representation hypothesis to a functional representation hypothesis, enabling concept learning beyond vector-valued representations. Unlike standard SAEs that employ multi-layer perceptrons (SAE-MLP) to each concept with a scalar activation, we introduce and formalize sparse autoencoder neural operators (SAE-NOs), which extend vector-valued representations to functional ones. We instantiate this framework as SAE Fourier neural operators (SAE-FNOs), parameterizing concepts as integral operators in the Fourier domain. We show that this functional parameterization fundamentally shapes learned concepts, leading to improved stability with respect to sparsity level, robustness to distribution shifts, and generalization across discretizations. We show that SAE-FNO is more efficient in concept utilization across data population and more effective in extracting localized patterns from data. We show that convolutional SAEs (SAE-CNNs) do not generalize their sparse representations to unseen input resolutions, whereas SAE-FNOs operate across resolutions and reliably recover the underlying representations. Our results demonstrate that moving from fixed-dimensional to functional representations extends sparse autoencoders from detectors of concept presence to models that capture the underlying structure of the data, highlighting parameterization as a central driver of interpretability and generalization.
@@ 1 Introduction
Sparse autoencoders (SAEs) (Makhzani and Frey, 2014) are popular in machine learning for learning structured and interpretable representations (Huben et al., 2023). More broadly, sparse representations are widely used for representation learning (Bengio et al., 2013), providing a mathematically grounded framework for uncovering parsimonious latent structure through interpretable components.
SAEs have increasingly become a tool for understanding model capabilities and internal representations themselves (Doshi-Velez and Kim, 2017). In this context, SAEs are widely used for mechanistic interpretability (Bricken et al., 2023a), where the goal is to explain model behaviour by identifying the internal concepts and representations that drive computation in large neural networks. The linear representation hypothesis (LRH) postulates that internal activations can be decomposed into orthogonal concepts. Hence, SAEs are primarily designed under this hypothesis to decompose activations into linearly accessible and approximately orthogonal concepts that can be sparsely and linearly combined to represent the data (Elhage et al., 2022).
LRH directly motivates the use of a shallow SAEs, where the learned decoder weights encapsulate the dictionary of interpretable concepts. This hypothesis has been supported by numerous empirical analyses on the geometry of concepts captured by internal representations of large models (Fel et al., 2023; Merullo et al., 2025). Given this hypothesis, the architecture of almost all SAEs for interpretability has shallow non-linear multi-layer perceptron (MLP) encoder and shallow linear decoder, as we term this as SAE-MLP, hence differing primarily on their choice on the encoder nonlinearity (Bussmann et al., 2025; Hindupur et al., 2025).
Figure 1: Sparse Autoencoder Neural Operator (SAE-NO). (a) Conventional SAEs (SAE-MLP) parameterize concepts ($\bm{D}$}) in the Euclidean domain, assigning each concept a scalar activation indicating presence and strength. (b) SAE-FNOs use functional concept parameterization via Fourier integral operators, encoding both concept sparsity and domain sparsity to capture how and where concepts are expressed. SAE-FNO treats inputs as samples from an underlying continuous function, learns concepts in the Fourier domain, and supports querying of representations, reconstructions, and concepts at varying resolutions.
While SAE-MLPs have proven effective for sparse model recovery (Olshausen and Field, 1996) and identifying monosemantic features (Bricken et al., 2023b), their architectural design inherently limits what a concept can represent. They all share three fundamental properties: i) concept utilization is summarized by a scalar activation, encoding presence and strength, but no further structure, ii) concepts are parameterized in the same domain as the inputs, and iii) the inference of sparse representation and the learning of concepts happen in Euclidean spaces with a fixed discretization sampling.
These properties introduce an inductive bias that can limit the ability of SAEs to learn generalized or localized concepts, or to be applied to data and activations across varying scales. In particular, scalar concept activations, especially when the underlying data exhibit spatial or temporal structure, cannot capture how a concept is expressed, where it manifests locally, in what orientation, or what structural properties of the input domain it encodes. Such capabilities are essential in scientific modelling, computer vision, and in analyzing model activations or physical data that exhibit locality or symmetric structure. While concept feature maps can be introduced through a convolutional parameterization of SAEs (which we introduce as SAE-CNNs), such models remain tied to a fixed resolution and inherit inductive biases imposed by parameterization in the input domain.
@@
This is in contrast with class of deep learning architectures known as neural operators (NO) (Li et al., 2021), which extends beyond Euclidean spaces and learns mappings between infinite-dimensional function spaces. Neural operators are central to data-driven scientific modelling (Azizzadenesheli et al., 2024), particularly for solving partial differential equations (PDEs), can generalize across discretizations, and have recently been applied in neuroscience to capture brain representations (Ghafourpour et al., 2025). While their predictive power is well established across applications (Li et al., 2024; Jatyani et al., 2025), their representational properties and capability for concept learning and mechanistic interpretability remain largely unknown.
Our contributions. We focus on extracting concepts from vision data, where local and spatial structure is inherent and where sparse autoencoders should capture such structure across both spatial locations and scales. We introduce sparse autoencoder neural operators (SAE-NOs), a new class of SAEs that operate directly in infinite-dimensional function spaces. We generalize the LRH to a functional representation hypothesis (FRH), enabling the learning and extraction of concepts in function spaces. Building on FRH, we instantiate SAE-NOs in the form of sparse autoencoder Fourier neural operators (SAE-FNOs).
With SAE-FNO, we introduce a principled tool for mechanistic interpretability (Nanda et al., 2023) in models and data with inherent locality, continuity, or spatial structure. Our results show that when such structure is essential to the underlying concepts, conventional SAEs fail to recover meaningful and expressive representations, while SAE-FNOs effectively capture and exploit these properties.
SAE-FNO parameterizes concepts as integral operators in the Fourier domain. We demonstrate that SAEs impose an inductive bias on the characteristics of learned concepts that depends critically on the domain in which the concepts are parameterized. To our knowledge, a systematic study of how concept parameterization shapes the inductive bias of SAEs has not been explored in prior work.
SAE-NOs replace scalar activations with feature-map representations, introducing domain sparsity alongside concept sparsity to encode where and how concepts are expressed. This offers structured representations that capture both concept presence and data structure. More precisely, concept sparsity acts as group sparsity (selecting active channels), while domain sparsity is intra-group spatial sparsity. While SAE-CNNs can exhibit a similar abstraction at a fixed resolution, they do not inherit the same benefits as SAE-FNOs in terms of generalization, efficiency of concept utilization, and stability across varying resolutions and sparsity levels.
We visualize the concepts in SAE-FNO and SAE-MLP and show that SAE-FNO exhibits superior expressivity, requiring fewer concepts. SAE-FNO learns concepts that are globally shared across the data population and are reused in different ways through the structure of their feature-map representations. As a result, SAE-FNO achieves significantly higher efficiency in concept utilization.
Compared to SAE-MLP and SAE-CNN, we show that SAE-FNO learns concepts with fundamentally different characteristics. In particular, SAE-FNO generalizes across resolutions, and its learned concepts are stable and robust to the choice of sparsity level and to distribution shifts. While many SAE variants exhibit an inductive bias toward learning approximately orthogonal concepts due to their shallow encoders, SAE-FNOs can learn concepts spanning a broader range of correlations even in the presence of this bias.
We introduce lifting into SAEs for concept extraction, mapping inputs to a higher-dimensional space where concept learning is performed. As a core component of neural operators, lifting increases expressivity and improves optimization. We show that lifting acts as a preconditioner that accelerates recovery and filters noise, enabling the learning of smoother and more robust concepts.
2 Preliminaries and Related Works
@@
The sparse linear generative model (Definition 2.1), also known as sparse coding (Olshausen and Field, 1996), has a rich history in statistics and in compressed sensing (Donoho, 2006; Candès et al., 2006). Representing data $\bm{x}$} under a sparse generative model involves learning a dictionary $\bm{D}$$^{*}$ to infer a sparse code $\bm{z}$}. We focus on the case where $\bm{D}$$^{*}$ is unknown (a.k.a. dictionary learning(Agarwal et al., 2016) and refer to it as sparse concept/model recovery, formulated as a bilevel problem (Definition 2.2(Tolooshams, 2023).
@@
A data sample ${\bm{x}}\in\mathbb{R}^{m}$ is said to follow a sparse generative model if there exists a sparse code ${\bm{z}}\in\mathbb{R}^{p}$ (with $\text{supp}({\bm{z}})\leq k\ll p$) and an overcomplete dictionary ${\bm{D}}^{*}\in\mathbb{R}^{m\times p}$ (with $p\gg m$) that model the data as ${\bm{x}}={\bm{D}}^{*}{\bm{z}}$ in the noiseless setting.
Definition 2.2 (Sparse Concept Recovery).
@@
Under the sparse generative model (Definition 2.1), the goal of sparse concept recovery is to estimate the underlying dictionary $\bm{D}$$^{*}$ by solving the following problem:
(outer) $\min_{{\bm{D}}\in\mathcal{D}}\quad\tfrac{1}{2}\ \|{\bm{x}}-{\bm{D}}{\bm{z}}\|_{2}^{2}+\lambda\mathcal{R}({\bm{z}})$ (1)
s.t.(inner) ${\bm{z}}=\text{Enc}_{\theta}({\bm{x}})=\operatorname*{arg\,min}_{{\bm{v}}\in\mathbb{R}^{p}}\mathcal{L}({\bm{x}},{\bm{v}},\theta),$
@@
where $\text{Enc}_{\theta}$ is a forward map encoder from data $\bm{x}$} to latent code $\bm{z}$}, defined as the minimizer of the inner loss $\mathcal{L}({\bm{x}},{\bm{v}},\theta)$. The outer-level optimizes $\bm{D}$}, and concept/model recovery is successful when the learned concept $\bm{D}$} approximates the true underlying dictionary $\bm{D}$$^{*}$.
@@
The connection between bilevel optimization and neural networks is well established through SAEs (Malézieux et al., 2022; Tolooshams and Ba, 2022). The inner and outer levels in Definition 2.2 mirror the encoder–decoder structure of a neural network, where the encoder $\text{Enc}_{\theta}$ maps data $\bm{x}$} to sparse codes $\bm{z}$}, and the decoder reconstructs $\bm{x}$} via ${\bm{D}}{\bm{z}}$ using learned concepts $\bm{D}$} (see Definition 2.3).
Definition 2.3 (Sparse Autoencoders for Concept Recovery).
@@
The goal of concept recovery with SAEs is to infer a sparse code ${\bm{z}}\in\mathbb{R}^{p}$, representing the data ${\bm{x}}\in\mathbb{R}^{m}$, with a linear combination of a dictionary of concepts $\bm{D}$}: $\quad\text{(encoder)}\quad{\bm{z}}=\text{Enc}_{\theta}({\bm{x}}),\quad\text{(decoder)}\quad\hat{\bm{x}}=\text{Dec}({\bm{z}})={\bm{D}}{\bm{z}}$, where $\theta$ denotes the set of encoder parameters. Model recovery refers to learning the concepts $\bm{D}$}.
The mechanistic interpretability community is interested in shallow autoencoders as $\text{Enc}({\bm{x}})=\mathcal{S}({\bm{W}}^{\top}{\bm{x}}+\bm{b})$. For mechanistic intepretability, the key distinguishing factor between SAEs is how they design the nonlinear activation (projection) $\mathcal{S}$ to be ReLU (Huben et al., 2023), TopK (Makhzani and Frey, 2014), SpaDE (Hindupur et al., 2025), JumpReLU (Rajamanoharan et al., 2024), BatchTopK (Bussmann et al., ), and MPSAE (Costa et al., 2025).
To extend SAEs to function spaces, we briefly review neural operators (Kovachki et al., 2023), which learn mappings between infinite-dimensional function spaces rather than finite-dimensional Euclidean spaces. Neural operators consist of three components: lifting, kernel integration, and projection (see Appendix). Lifting maps data to a higher-dimensional channel space where learning occurs, followed by projection back to a lower-dimensional space, and improves the optimization of SAEs.
@@
Fourier Neural Operator (FNO) (Li et al., 2021) models the kernel integral operator (Definition A.2) with a convolution operator parameterized in Fourier space (Definition A.3Definition A.4). The convolution operator parameterizes the kernel $\kappa(x,y)=\kappa(x-y)$ as a complex function $\kappa:D\rightarrow\mathbb{C}^{d_{w}\times d_{v}}$. FNO is used with truncated frequency modes in practice, shown to improve performance and lower sensitivity to decreasing the discretization sampling.
3 Sparse Autoencoder Neural Operators
We extend SAEs, typically implemented with MLP-based encoders (SAE-MLPs), to neural operators to enable concept recovery in function spaces. While standard SAEs represent each concept with a scalar activation indicating its presence and strength, we instead learn global feature-map representations whose energy corresponds to the degree to which a concept is utilized. Importantly, the representations encode both localized and global structure, capturing how each concept contributes to representing the data (Figure 1).
To introduce SAE-NOs, we define a sparse functional generative model of data to formalize the functional representation hypothesis (FRH) (Definition 3.1) and formulate its corresponding optimization problem for concept recovery in function spaces (Definition 3.2).
Definition 3.1 (Functional Representation Hypothesis (FRH)).
@@ 3 Sparse Autoencoder Neural Operators
Consider a linear operator $\mathcal{G}_{{\bm{D}}^{*}}:\mathcal{Z}\rightarrow\mathcal{X}$ in function spaces, where $\mathcal{Z}$ and $\mathcal{X}$ are Banach spaces of functions defined on bounded domains $D_{z}\subset\mathbb{R}^{p}$ and $D_{x}\subset\mathbb{R}^{m}$, respectively. The data sample ${\bm{x}}\in\mathcal{X}$ is said to follow a sparse functional generative model or the functional representation hypothesis (FRH) if there exists a sparse latent representation $\bm{z}$} drawn from the support $\mathcal{Z}$ of a probability measure $\mu$, where ${\bm{x}}=\mathcal{G}_{{\bm{D}}^{*}}({\bm{z}})$.
Definition 3.2 (Sparse Functional Concept Recovery).
@@
Under the sparse functional generative model (Definition 3.1), the goal of sparse concept recovery in function spaces is to estimate the underlying operator $\mathcal{G}_{{\bm{D}}^{*}}$ by solving the following bilevel optimization problem:
$\min_{{\bm{D}}\in\mathcal{D}}$ $\mathbb{E}_{{\bm{z}}\sim\mu}\|\mathcal{G}_{{\bm{D}}^{*}}({\bm{z}})-\mathcal{G}_{{\bm{D}}}(\hat{\bm{z}})\|_{\mathcal{X}}^{2}$ (2)
s.t. $\hat{\bm{z}}=\operatorname*{arg\,min}_{{\bm{v}}}\ \mathbb{E}_{{\bm{z}}\sim\mu}\|{\bm{v}}-\mathcal{H}_{\theta}(\mathcal{G}_{{\bm{D}}^{*}}({\bm{z}}))\|_{\mathcal{Z}}^{2},$
@@
where the operator $\mathcal{H}_{\theta}:\mathcal{X}\rightarrow\mathcal{Z}$ aims to approximate the inverse of $\mathcal{G}_{{\bm{D}}^{*}}$. Model recovery is successful when the learned concepts $\bm{D}$} in $\mathcal{G}_{{\bm{D}}}$ closely approximate $\bm{D}$$^{*}$ in $\mathcal{G}_{{\bm{D}}^{*}}$.
Figure 2: Concept Recovery Examined Under Known Generative Model. (a) SAE-FNOs can recover underlying concepts. The dictionary (concept) recovery error converges to zero as a function of training iterations. (b) Lifting the SAE onto a high-dimensional space for concept learning acts as a preconditioner and accelerates recovery. (c) SAE-FNO is superior to SAE-CNN in recovering smooth concepts. Each dot represents one training scenario with x-axis as the initialized concept error and y-axis as the final trained concept error.
(a) SAE-FNO (original rate).
(b) SAE-FNO ($4\times$ upsampling rate).
(c) SAE-CNN (original rate).
(d) SAE-CNN ($4\times$ upsampling rate).
Figure 3: Discretization Generalization of SAE-FNO. The underlying data structure has five convolutional concepts along with a 55-sparse code (i.e., 11-sparse representation for each concept). Each subplot shows (left) the input data and its reconstruction, and (right) the support sign (active sparse codes). (a-b) SAE-FNO can successfully infer the underlying sparse representations and reconstruct the data at the original as well as an increased rate. (c-d) While SAE-CNN can reconstruct the data and infer the sparse representation at the original rate, it fails to represent the data at a resolution different from the one trained at (see also Figure 11).
@@
To recover concepts, we formalize an SAE-NO (Definition 3.3) with an encoder approximating the inverse generative operator and a decoder that captures the underlying operator and its learnable concepts.
Definition 3.3 (Sparse Autoencoder Neural Operators (SAE-NOs) for Concept Recovery).